5. What are E cell and AG of a redox reaction at 25 °C for which n= 1 and K = 5.0 x 104?
Answer #1

\DeltaGo = - 2.303 RT log K ; where R is universal gas constant, T is absolute temperature and K is the equilibrium constant.

Given T= 298 K (250 C)

K = 5.0*104 and we know

R= 8.314 J/mol.K

\DeltaGo = -2.303* 8.314 J/mol.K *298 K* log (5*104) = -26811.61 J/mol

Again \DeltaGo = -n.F.Eocell  where n= 1 (given), F= 96500 C,  \DeltaGo = -26811.61 J (calculated)

Eocell = -(\DeltaGo/ nF) = -(-26811.61/ 1*96500) = 0.278 J/C = 0.278 V

If you still have any confusion feel free to ask.

Do rate if you got it useful.

Thanks,